OpenStudy (anonymous):

When 4 inches are added to each side of square, its area increases by 64 square inches. What is the area of the initial square?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

\[A=x ^{2}\] with x being the length of the initial sides \[A+64=(x+4)\]

OpenStudy (anonymous):

the area of previous square was 36 sq. inches

OpenStudy (anonymous):

\[(x+4)^{2}\] sorry

OpenStudy (anonymous):

(x + 4)^2 = x^2 + 64. This will be a quadratic in "x".

OpenStudy (anonymous):

how did you get 36 so i can do the rest by myself?

OpenStudy (anonymous):

it can be done by trial and error also

OpenStudy (anonymous):

(x + 4)^2 = x^2 + 64 will become x^2 + 8x + 16 = x^2 + 64. Just combine terms and simplify.

OpenStudy (anonymous):

ok one thing why do you square it

OpenStudy (anonymous):

area of square is its side squared

OpenStudy (anonymous):

You have to get the area. Both of the first square and the second.

OpenStudy (anonymous):

and x is the side

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

The second area is (x + 4)^2 and the first area is x^2

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

its our pleasure

OpenStudy (anonymous):

\[A _{1}=x _{1}^{2}\] \[A _{2}=x _{2}^{2}\] \[A _{2}=A _{1}+64\] \[x _{2}=x _{1}+4\] Does that help?

OpenStudy (anonymous):

lol i got it

OpenStudy (anonymous):

ok so i did x^2 + 8x + 16 = x^2 + 64. and i got 6 what did i do wrong

OpenStudy (anonymous):

you got \[x _{1}=6\] so what's the area? Never forget what was the initial question